# CompBench: Benchmarking Complex Instruction-guided Image Editing

### v2

@@ 3.2 Dataset Generation
In this section, we detailedly demonstrate the generation process of our CompBench. The overall pipeline is shown in Figure 3.
Source Data Collection and Preprocessing. To address the scarcity of high-quality paired complex editing data, we select MOSE [5], a video instance segmentation (VOS) dataset featuring complex scenes with multi-object masks. The dataset undergoes a rigorous preprocessing pipeline: We first filter low-quality video frames using a mixture of no-reference image quality assessment metrics (e.g., NIQE [42]) to eliminate blurry, low-contrast, or corrupted samples. Then, A professional team manually verifies the filtered data, retaining only high-quality images. For mask preprocessing, multi-object masks are decomposed into single-object masks to isolate editable entities. A multimodal large language model (MLLM, e.g., Qwen-VL [34]) evaluates mask continuity and occlusion levels, discarding discontinuous and incomplete masks with high occlusion rate. Similarly, annotators further check these masks to ensure pixel-level precision.
@@
Task-specific Data Generation Pipelines. To address the unique challenges and diversity of complex instruction-guided image editing tasks, we design four specialized data construction pipelines tailored to distinct task categories: (1) local editing pipeline for object-level manipulations (object removal, object addition, object replacement). (2) action/scene spatial editing pipeline for modifying object dynamics or scene perspectives (action editing, location editing, viewpoint editing). (3) complex reasoning pipeline for implicit contextual edits requiring reasoning (implicit reasoning). (4) multi-editing pipeline for multi-object and multi-turn editing tasks. All pipelines adopt a unified MLLM-Human Collaborative Framework: multimodal large language models (MLLMs) [18, 1, 20, 34] generate initial task-specific instructions by analyzing visual scenes and editing goals, followed by human validation to ensure instruction-image semantic alignment and image editing fidelity. Unsuccessful edits are iteratively re-generated or discarded, retaining only high-fidelity samples that satisfy both linguistic precision and visual realism. During inference, models directly process free-form natural language instructions without pipeline-specific constraints, eliminating intermediate processing steps and enhancing operational efficiency. Detailed implementation procedures for each pipeline are provided in the appendix A.
@@
Figure 4: Characteristics and statistics of CompBench. (a) Task taxonomy of CompBench, illustrating the full range of task types. (b) SSIM [35] comparison among different datasets and benchmarks. Note that UltraEdit [45] and InstructPix2pix [2] are datasets, whereas the remaining entries are benchmarks.
@@
Instruction Decomposition Strategy. To improve the construction of image editing instructions with greater clarity and precision, we propose a structured framework that organizes editing instructions along four complementary aspects: spatial positioning, visual attributes, motion states, and object entities. This structured approach transforms potentially ambiguous editing requirements into well-defined specifications while maintaining natural expressiveness. The method employs a two-phase generation process: first, multimodal large language models produce dimension-aware instruction candidates by analyzing visual contexts. Then human experts refine these to ensure precision and consistency. By systematically addressing each aspect of the editing operation while preserving the flexibility of natural language, this framework enables the creation of instructions that are both intuitively understandable and technically precise for complex image editing tasks.
Characteristics and Statistics. As illustrated in Figure 4(a), our benchmark comprises 5 major categories encompassing a total of 9 complex editing tasks, yielding 3k+ image editing samples with corresponding complex instructions. Since we lack global captions for pre-edited and post-edited images, we employ the Structural Similarity Index Measure (SSIM) [35] to evaluate the semantic consistency between image pairs as a quality assessment metric. As shown in Figure 4(b), CompBench achieves notably higher SSIM than other datasets and benchmarks.
@@
Notably, our dataset features significantly more challenging editing tasks, each requiring comprehensive capabilities such as visual grounding and complex reasoning. Detailed analysis of core competencies essential for our benchmark are discussed in Appendix B. To systematically evaluate scene complexity, we adopt multiple quantitative indicators such as average number of scene objects and categories. These metrics demonstrate that our benchmark exhibits substantially higher complexity compared to existing benchmarks.
@@ 4.1 Settings
Baselines. Given that our study specifically targets instruction-guided image editing tasks, we restrict our selection to instruction-guided image editing models and exclude approaches based on global description guidance. The evaluated models include: InstructPix2pix [2], MagicBrush [38], HIVE [44], Smart-edit [14], MGIE [9], HQ-Edit [15], CosXL-Edit [30], UltraEdit [45], AnyEdit [41], Seed-X-Edit [10], GoT [7], and the recently released Step1X-Edit [21].
@@
Model Local Editing Multi-object Editing Implicit Reasoning Foreground Background Foreground Background Foreground Background LC-T $\uparrow$ LC-I $\uparrow$ PSNR(dB) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ LC-T $\uparrow$ LC-I $\uparrow$ PSNR(dB) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ LC-T $\uparrow$ LC-I $\uparrow$ PSNR(dB) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ InstructPix2pix [2] 19.445 0.777 21.416 0.695 0.137 19.624 0.779 20.200 0.658 0.162 19.007 0.793 21.806 0.683 0.125 MagicBrush [38] 20.067 0.798 23.415 0.744 0.088 19.878 0.800 23.524 0.727 0.092 19.526 0.828 22.145 0.714 0.106 HIVE-w [44] 19.820 0.770 19.908 0.641 0.198 20.024 0.775 19.594 0.609 0.224 18.634 0.777 20.268 0.602 0.219 HIVE-c [44] 19.236 0.772 21.741 0.689 0.147 19.585 0.781 21.560 0.663 0.155 18.938 0.786 22.168 0.666 0.132 Smart-edit-7B [14] 20.034 0.798 24.398 0.761 0.073 19.888 0.803 23.959 0.731 0.090 19.743 0.832 23.058 0.732 0.096 MGIE [9] 18.957 0.780 20.572 0.708 0.175 19.830 0.788 18.296 0.692 0.274 17.728 0.801 24.432 0.780 0.088 CosXL-Edit [30] 19.029 0.778 20.442 0.706 0.156 19.550 0.788 20.382 0.682 0.171 18.269 0.794 20.984 0.681 0.161 HQ-Edit [15] 18.316 0.734 12.240 0.419 0.441 19.163 0.757 12.987 0.412 0.421 18.864 0.767 12.321 0.396 0.452 UltraEdit [45] 19.618 0.786 22.938 0.783 0.145 20.022 0.795 22.326 0.719 0.164 18.350 0.784 23.374 0.717 0.145 AnyEdit [41] 19.932 0.794 22.769 0.714 0.125 19.875 0.809 22.789 0.697 0.129 19.588 0.816 20.271 0.639 0.191 SEED-X [10] 17.933 0.780 21.466 0.805 0.139 19.092 0.795 20.638 0.788 0.158 17.467 0.784 21.506 0.709 0.134 GoT [7] 20.268 0.807 24.675 0.890 0.067 19.919 0.804 21.296 0.826 0.127 19.237 0.820 24.738 0.860 0.088 Step1X-Edit [21] 20.501 0.817 23.371 0.882 0.078 20.213 0.828 22.696 0.873 0.089 19.312 0.850 23.435 0.869 0.082
@@
Model Turn1 Turn2 Foreground Background Foreground Background LC-T LC-I PSNR SSIM LPIPS LC-T LC-I PSNR SSIM LPIPS InstructPix2pix [2] 19.424 0.784 21.073 0.676 0.142 19.818 0.776 17.607 0.568 0.238 MagicBrush [38] 19.977 0.812 24.020 0.730 0.089 20.253 0.811 21.244 0.682 0.134 HIVE-w [44] 19.784 0.781 20.040 0.621 0.196 20.129 0.761 17.291 0.532 0.272 HIVE-c [44] 19.756 0.787 21.330 0.660 0.155 19.812 0.778 18.346 0.590 0.217 Smart-edit-7B [14] 19.876 0.817 24.632 0.740 0.080 20.050 0.807 23.404 0.724 0.104 MGIE [9] 19.355 0.801 21.563 0.731 0.143 19.695 0.798 18.382 0.655 0.223 HQ-Edit [15] 18.987 0.755 12.950 0.410 0.422 18.935 0.740 12.032 0.383 0.499 CosXL-Edit [30] 19.389 0.787 20.233 0.679 0.171 19.394 0.771 16.752 0.590 0.301 UltraEdit [45] 19.990 0.792 23.763 0.715 0.116 20.177 0.779 22.917 0.715 0.139 AnyEdit [41] 19.953 0.812 23.412 0.711 0.113 20.093 0.803 20.010 0.633 0.188 SEED-X [10] 19.139 0.795 21.042 0.792 0.153 18.729 0.753 13.793 0.457 0.404 GoT [7] 20.108 0.816 25.089 0.894 0.066 19.939 0.804 21.397 0.825 0.131 Step1X-Edit [21] 20.157 0.832 23.987 0.883 0.078 20.262 0.835 20.710 0.822 0.128
Evaluation Metrics and Methods. Evaluation metrics for image editing tasks must be well-suited to the complexity of our scenarios, providing a comprehensive and accurate assessment of editing performance in complex scenes. Moreover, the metrics should be tailored to reflect the unique characteristics of different task types. For tasks including local editing, multi-editing and implicit reasoning, we posit that an effective editing model should modify foreground objects while preserving background consistency. Therefore, we evaluate the editing performance from both foreground and background perspectives. For background consistency assessment, we compute three metrics, including PSNR, SSIM [35], and LPIPS [43], on the background regions. For the foreground evaluation, we consider two aspects: editing accuracy and instruction following. For editing accuracy, we measure the similarity between the edited result and the ground truth (GT) image by comparing their CLIP [24] image embeddings in the foreground region, thereby determining whether the edited foreground visually aligns with the GT. To assess instruction-following capability, we measure the CLIP [24] similarity between the edited foreground object and the textual description of the target region to evaluate the model’s ability to interpret and execute the given instructions. Additionally, for action editing, location editing and viewpoint editing tasks, where the object’s morphology, position, or viewpoint may change significantly, the aforementioned metrics are insufficient for evaluation. Thus, we employ GPT-4o [23] for assessment. Specifically, we design tailored prompts for these three tasks, instructing GPT-4o to evaluate the editing performance from multiple perspectives and assign a score between 0 and 10. The detailed prompt designs are provided in the Appendix D.
@@
Table 4: Evaluation results on action editing, location editing and viewpoint editing tasks.
@@
Model Action Location Viewpoint Average InstructPix2pix [2] 3.047 3.425 0.088 2.187 MagicBrush [38] 3.511 4.603 0.500 2.871 HIVE-w [44] 3.151 4.110 0.559 2.607 HIVE-c [44] 3.977 4.192 0.559 2.909 Smart-edit-7B [14] 4.233 3.890 1.412 3.178 MGIE [9] 1.921 1.726 0.059 1.235 HQ-Edit [15] 1.449 1.425 0 0.958 CosXL-Edit [30] 4.270 5.479 0.235 3.328 UltraEdit [45] 4.449 4.014 0.235 2.899 AnyEdit [41] 3.750 5.068 0.471 3.096 SEED-X [10] 2.270 3.028 1.412 2.237 GoT [7] 3.337 3.625 0.206 2.389 Step1X-Edit [21] 6.270 5.041 1.500 4.270
@@ 4.2 Experiment Results
The experimental results for local editing, multi-turn editing, multi-object editing, implicit reasoning, and action/location/viewpoint editing are presented in Tables 2,  3, and  4, respectively. Our key analysis of the results are as follows: (1) No model dominates across all tasks. Among all evaluated models, Step1X-Edit [21] emerges as the most prominent one, achieving top results in 17 out of 29 metrics (nearly 60%) across 9 tasks. Notably, Step1X-Edit [21], SmartEdit [14], and GoT [7] consistently deliver superior performance, securing top-three rankings in the majority of metrics across most tasks. In contrast, HQ-Edit [15] demonstrates substantially inferior results in nearly all tasks. (2) For multi-turn editing tasks, all models exhibit a notable decline in background consistency metrics during the second editing round. Among them, SmartEdit [14] maintains relatively robust performance in second editing turn. In multi-object editing, Step1X-Edit [21] outperforms other models by a significant margin, showcasing exceptional capability in simultaneous multi-object manipulation. (3) Beyond the consistently high-performing models, MGIE [9] also demonstrates competitive results in background consistency for implicit reasoning tasks. (4) For action/location/viewpoint editing tasks, GoT [7] does not perform as strongly compared to its results in other tasks. Step1X-Edit [21] remains the most reliable performer, achieving the only score above 6 (out of 10) in the highly challenging action editing task. Meanwhile, Smartedit [14] and CosXL-Edit [30] exhibit strong performance in these tasks, distinguishing themselves from other models.

### v4

@@ 3.2 Dataset Generation
In this section, we detailedly demonstrate the generation process of our CompBench(Figure 3).
Source Data Collection and Preprocessing.
We utilize the MOSE dataset [7] to address the lack of complex editing data. Image quality is first guaranteed by filtering corrupted frames via automated metrics (e.g., NIQE [45]) and subsequent manual verification. For masks, we decompose multi-object annotations into single-object instances. Finally, discontinuous or heavily occluded masks are discarded using MLLMs (e.g., Qwen-VL [35]), followed by manual refinement to ensure pixel-level precision.
Task-specific Data Generation Pipelines.
@@
We design four specialized pipelines to cover diverse complex instruction-guided image editing scenarios: (1) local editing pipeline for object-level manipulations (object removal, object addition, object replacement). (2) action/scene spatial editing pipeline for modifying object dynamics or scene perspectives (action editing, location editing, viewpoint editing). (3) complex reasoning pipeline for implicit contextual edits requiring reasoning (implicit reasoning). (4) multi-editing pipeline for multi-object and multi-turn editing tasks. All pipelines adopt a unified MLLM-Human Collaborative Framework: multimodal large language models (MLLMs) [18, 1, 20, 35] generate initial task-specific instructions by analyzing visual scenes and editing goals, followed by human validation to ensure instruction-image alignment and image editing fidelity. Unsuccessful edits are iteratively re-generated or discarded, retaining only high-fidelity samples. Detailed procedures are provided in the supplementary material.
Figure 4: Task taxonomy of CompBench. Illustration of the full range of task types in CompBench.
@@
Figure 5: SSIM comparison among different datasets and benchmarks. Note that UltraEdit [48] and InstructPix2pix [3] are datasets, whereas the remaining entries are benchmarks.
Instruction Decomposition Strategy.
@@
To enhance the clarity and precision of editing instructions, we propose a structured framework that organizes editing instructions along four aspects: spatial positioning, visual attributes, motion states, and object entities. This approach transforms ambiguous editing requests into well-defined specifications while maintaining natural expressiveness. The method employs a two-phase generation process: first, an MLLM produces dimension-aware instruction candidates by analyzing visual contexts. Then human experts refine these to ensure precision and consistency. By systematically addressing each aspect of the editing operation while preserving the flexibility of natural language, this framework enables the creation of instructions that are both intuitively understandable and technically precise for complex image editing tasks.
Characteristics and Statistics.
@@
As illustrated in Figure 4, our benchmark comprises five major categories encompassing nine complex editing tasks, yielding over 3000 high-quality samples. We employ SSIM [36] to evaluate semantic consistency between pre- and post-edited images. As shown in Figure 5, CompBench achieves notably higher SSIM than other datasets and benchmarks. Notably, our dataset features significantly more challenging editing tasks requiring comprehensive capabilities such as visual grounding and complex reasoning. Quantitative indicators (e.g., average number of objects and categories) also demonstrate that our benchmark features substantially higher scene complexity. Detailed subtask definitions and core competency analyses are provided in the supplementary material.
@@ 4.1 Settings
Baselines. Our study focuses on instruction-guided image editing models, excluding those based on global description guidance. The evaluated models include: InstructPix2pix [3], MagicBrush [44], HIVE [47], Smart-edit [14], MGIE [11], HQ-Edit [15], CosXL-Edit [31], UltraEdit [48], AnyEdit [42], Seed-X-Edit [12], GoT [9], Step1X-Edit [21], Bagel [5], FLUX.1 Kontext [16], and Qwen-Image-Edit [38].
@@
Model Local Editing Multi-object Editing Implicit Reasoning Foreground Background Foreground Background Foreground Background LC-T $\uparrow$ LC-I $\uparrow$ PSNR(dB) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ LC-T $\uparrow$ LC-I $\uparrow$ PSNR(dB) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ LC-T $\uparrow$ LC-I $\uparrow$ PSNR(dB) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ InstructPix2pix [3] 19.269 0.778 21.828 0.706 0.124 20.050 0.804 20.534 0.671 0.152 18.981 0.794 21.813 0.683 0.125 MagicBrush [44] 20.051 0.798 23.429 0.741 0.090 20.004 0.821 24.176 0.738 0.081 19.498 0.826 22.143 0.714 0.106 HIVE-w [47] 19.804 0.771 19.903 0.641 0.198 19.949 0.782 19.896 0.624 0.210 18.590 0.777 20.261 0.602 0.219 HIVE-c [47] 19.226 0.773 21.732 0.689 0.147 19.904 0.798 21.854 0.676 0.144 18.888 0.787 22.167 0.666 0.144 Smart-edit-7B [14] 19.999 0.799 24.389 0.761 0.074 20.186 0.825 24.886 0.745 0.076 19.740 0.831 23.060 0.732 0.096 MGIE [11] 18.773 0.764 18.204 0.639 0.257 20.102 0.742 15.380 0.500 0.354 18.304 0.795 22.144 0.719 0.151 CosXL-Edit [31] 19.068 0.778 20.809 0.712 0.148 19.606 0.807 21.068 0.698 0.153 18.003 0.800 21.190 0.683 0.156 HQ-Edit [15] 18.888 0.734 11.768 0.411 0.428 19.401 0.708 13.032 0.432 0.400 18.688 0.766 11.877 0.387 0.459 UltraEdit [48] 19.690 0.787 22.932 0.741 0.143 20.081 0.810 22.725 0.732 0.152 18.289 0.786 23.362 0.717 0.142 AnyEdit [42] 19.994 0.797 22.812 0.716 0.124 20.257 0.810 23.493 0.710 0.116 19.572 0.816 20.276 0.639 0.191 SEED-X [12] 17.900 0.780 21.456 0.805 0.139 19.418 0.835 21.166 0.798 0.148 17.437 0.782 21.438 0.790 0.134 GoT [9] 20.268 0.807 24.675 0.889 0.067 20.225 0.827 22.486 0.842 0.108 19.246 0.821 24.889 0.860 0.088 Step1X-Edit [21] 20.495 0.817 23.372 0.882 0.078 20.459 0.861 23.782 0.886 0.078 19.175 0.847 23.408 0.869 0.083 Bagel [5] 21.059 0.838 27.692 0.935 0.045 20.856 0.883 26.849 0.932 0.051 19.719 0.874 28.891 0.919 0.052 FLUX.1 Kontext [16] 21.328 0.821 25.613 0.941 0.050 21.020 0.868 26.278 0.952 0.048 19.596 0.867 25.401 0.932 0.061 Qwen-Image-Edit [38] 21.522 0.828 24.968 0.892 0.072 20.722 0.864 23.492 0.826 0.103 20.046 0.859 22.815 0.775 0.124
@@
Model Turn1 Turn2 Foreground Background Foreground Background LC-T LC-I PSNR SSIM LPIPS LC-T LC-I PSNR SSIM LPIPS InstructPix2pix [3] 19.617 0.786 21.361 0.682 0.134 19.856 0.777 17.764 0.573 0.233 MagicBrush [44] 19.564 0.811 24.099 0.731 0.089 19.694 0.809 21.294 0.683 0.133 HIVE-w [47] 20.012 0.781 20.065 0.622 0.195 20.389 0.773 17.315 0.533 0.270 HIVE-c [47] 19.922 0.789 21.331 0.660 0.154 19.964 0.779 18.350 0.590 0.217 Smart-edit-7B [14] 19.595 0.813 24.653 0.740 0.080 19.571 0.808 23.383 0.723 0.104 MGIE [11] 19.194 0.814 18.473 0.602 0.246 19.328 0.803 14.820 0.493 0.366 HQ-Edit [15] 19.364 0.758 12.384 0.391 0.410 19.228 0.753 11.608 0.351 0.478 CosXL-Edit [31] 19.670 0.790 20.292 0.681 0.167 19.430 0.769 16.390 0.566 0.312 UltraEdit [48] 19.816 0.793 23.065 0.737 0.151 20.015 0.793 19.927 0.666 0.208 AnyEdit [42] 19.858 0.812 23.398 0.711 0.113 20.061 0.806 20.001 0.633 0.189 SEED-X [12] 19.576 0.796 21.049 0.792 0.153 19.188 0.769 17.708 0.629 0.280 GoT [9] 19.833 0.815 24.840 0.891 0.067 19.832 0.809 23.487 0.855 0.107 Step1X-Edit [21] 19.873 0.833 23.802 0.888 0.078 19.754 0.834 21.016 0.832 0.125 Bagel [5] 20.060 0.852 28.959 0.949 0.036 20.007 0.850 23.941 0.897 0.085 FLUX.1 Kontext [16] 20.634 0.845 26.193 0.954 0.046 20.620 0.840 22.417 0.907 0.094 Qwen-Image-Edit [38] 20.468 0.845 24.863 0.846 0.088 20.584 0.838 20.814 0.783 0.151
Evaluation Metrics and Methods.
For local editing, multi-editing, and implicit reasoning tasks, we employ a foreground-background decoupling strategy. To evaluate background preservation, we compute PSNR, SSIM [36], and LPIPS [46] on the background regions. For foreground evaluation, we assess two aspects using CLIP [25]: editing accuracy via CLIP image similarity between the localized edited foreground and the ground truth (GT) image (denoted as LC-I), and instruction following via local CLIP scores between the edited foreground and the local description (denoted as LC-T).
Additionally, for action, location, and viewpoint editing tasks—where the object’s morphology, position, or viewpoint may change significantly—automatic metrics alone are insufficient. To address this, we introduce multi-perspective scoring via GPT-4o [23], Qwen2.5-VL-72B [2], and human annotators. We design tailored prompts instructing the models to rate editing performance on a 0-10 scale. Simultaneously, trained human annotators evaluate background fidelity, editing intent, instruction following, and artifact presence based on standardized guidelines. Detailed prompts, annotation instructions, and additional evaluation results are available in the supplementary material.
@@
Table 4: Comparison on Action, Location, and Viewpoint Editing. Results for GPT-4o, Qwen-72B, Human Evaluation, and Average scores (top-3 per column highlighted in red, blue, green).
@@
Model Action Location Viewpoint GPT Qwen Human Avg. GPT Qwen Human Avg. GPT Qwen Human Avg. InstructPix2pix [3] 3.047 1.124 3.101 2.424 3.425 2.167 2.859 2.817 0.699 0.482 0.036 0.406 MagicBrush [44] 3.511 1.449 3.584 2.848 4.603 2.260 3.717 3.527 0.892 0.410 0.108 0.470 HIVE-w [47] 3.151 1.764 3.067 2.661 4.110 2.192 3.421 3.241 1.494 0.283 0.036 0.604 HIVE-c [47] 3.977 1.596 3.797 3.123 4.192 2.470 3.558 3.407 2.193 0.675 0.145 1.004 Smart-edit-7B [14] 4.233 1.607 4.348 3.396 3.890 2.875 3.505 3.423 2.169 0.590 0.410 1.056 MGIE [11] 1.921 1.213 1.797 1.644 1.726 1.795 1.728 1.750 0.205 0.193 0 0.133 CosXL-Edit [31] 4.270 2.375 3.966 3.537 5.479 2.493 4.517 4.163 1.916 0.988 0.301 1.068 HQ-Edit [15] 1.449 0.528 1.033 1.003 1.425 0.726 1.079 1.077 0.470 0.289 0 0.253 UltraEdit [48] 4.449 1.807 4.235 3.497 4.014 2.055 3.410 3.160 0.494 0.706 0 0.400 AnyEdit [42] 3.750 0.978 3.168 2.632 5.068 2.479 4.178 3.908 1.687 0.783 0.072 0.847 SEED-X [12] 2.270 1.494 1.685 1.816 3.028 3.247 2.771 3.015 2.241 1.169 0 1.137 GoT [9] 3.337 1.989 3.134 2.820 3.625 3.192 3.164 3.327 0.916 0.675 0.446 0.679 Step1X-Edit [21] 6.270 3.944 5.348 5.187 5.041 4.479 4.786 4.769 2.470 1.205 0.663 1.446 Bagel [5] 6.899 5.056 6.629 6.195 7.137 6.233 6.219 6.530 5.193 3.892 4.663 4.583 FLUX.1 Kontext [16] 5.169 3.202 4.517 4.296 3.000 3.110 3.836 3.315 3.471 2.373 3.108 2.984 Qwen-Image-Edit [38] 6.910 5.382 6.764 6.352 7.055 5.096 4.658 5.603 6.193 4.470 6.181 5.615
@@ 4.2 Experiment Results
The experimental results for local editing, multi-turn editing, multi-object editing, implicit reasoning, and action/location/viewpoint editing are presented in Tables 2,  3, and  4, respectively. Our key analysis of the results are as follows: (1) No model dominates across all tasks. Among all evaluated models, Bagel [5] emerges as the most prominent one, achieving top results in 18 out of 37 metrics (nearly 50%) across 9 tasks. Notably, Bagel [5], Qwen-Image-Edit [38], and FLUX.1 Kontext [16] consistently deliver superior performance, securing top-three rankings in the majority of metrics across most tasks, followed by Step1X-Edit [21]. In contrast, HQ-Edit [15] demonstrates substantially inferior results in nearly all tasks. (2) For multi-turn editing tasks, all models exhibit a notable decline in background consistency metrics during the second editing round. Among them, SmartEdit [14] maintains relatively robust performance in the second editing turn. (3) Bagel [5] consistently leads in the LC-I metric, reflecting its superior foreground fidelity. Additionally, it ranks high in background consistency, effectively preserving spatial and contextual information during editing. (4) For the more challenging action/location/viewpoint editing tasks, Qwen-Image-Edit [38] and Bagel [5] perform comparably and significantly outperform most other models. Step1X-Edit [21] also exhibits promising editing performance in these scenarios.
Figure 6: Overall Model Performance.(a) Top 5 model performance in five major tasks. (b) Overall model performance across all tasks.